OpenStudy (anonymous):

systems of equations : y=x-3 4x+y=32 how do you solve this?

OpenStudy (anonymous):

You know that y=x-3. You can use this in y our second equation to solve for x 4x+(x-3)=32 5x-3=32 5x=35 x=7 Now you know x, so you can substitute 7 for x in y=x-3. y=7-3=4 You can check if it's correct by plugging both values back into the other equation: 4(7)+4=32=28+4=32. The expression is true, so those are the solutions.

OpenStudy (anonymous):

wait is y already solved?

OpenStudy (anonymous):

No, y is not already solved. I had to solve it after I found x here. We substituted x-3 for y in the other equation so we could solve for x.

OpenStudy (anonymous):

\[4x+y=4x+(x-3)=32\]